The following contingency table shows the number of customers who bought various brands of shoes at Amazon, Outlets and Macys (No Excel)

Brand

Amazon

Outlets

Macys

Nike

48

35

37

Adidas

36

24

30

UGG

25

12

44

Given that the shoes was purchased at Macys, the probability that it is Nike is______

(No Excel)
Answer #1

Consider the below table.

Brand Amazon Outlets Macys Total
Nike 48 35 37 120
Adidas 36 24 30 90
UGG 25 12 44 81
Total 109 71 111 291

P(Nike | Macys) = P(Nike Macys) / P(Macys)

                          =(37/291) / (111/291) .......(from table and P(A) = number of occurance of event A divided by the number of occurance of all events)

                          = 37 / 111

                          = 1 / 3 = 0.3333

The probability that it is Nike given that the shoes was purchased at Macys is 0.3333.